# Magic tricycles: efficient magic state generation with finite block-length quantum LDPC codes

### v1

@@ Abstract
The preparation of high-fidelity non-Clifford (magic) states is an essential subroutine for universal quantum computation, but imposes substantial space-time overhead. Magic state factories based on high rate and distance quantum low-density parity check (LDPC) codes equipped with transversal non-Clifford gates can potentially reduce these overheads significantly, by circumventing the need for multiple rounds of distillation and producing a large number of magic states in a single code-block. As a step towards realizing efficient, fault-tolerant magic state distillation, we introduce a class of finite block-length quantum LDPC codes which we name tricycle codes, generalizing the well-known bicycle codes to three homological dimensions. These codes can support constant-depth physical circuits that implement logical CCZ gates between three code blocks. We show that these tricycle codes enable magic state generation with single-shot state-preparation and error correction, leading to a deterministic low-overhead distillation protocol without requiring post-selection. Numerical simulations of specific codes confirm robust performance under circuit-level noise with a Belief-Propagation+Order-Statistics Decoder (BPOSD), demonstrating a high circuit-noise threshold of $0.4\%$. Finally, we construct optimal depth syndrome extraction circuits for tricycle codes and present a protocol for implementing them efficiently on a reconfigurable neutral atom array platform.
@@ 1.1 Background
Over the last decade, remarkable progress has been made towards significantly reducing the overhead associated with magic state distillation by designing various codes with transversal non-Clifford gates and favorable code parameters [14, 50, 40, 21]. In fact, it has very recently been shown that families of quantum codes with transversal non-Clifford gates and asymptotically constant rate and relative distance exist, leading in principle to constant spatial overhead MSD protocols [63, 94]. However, these proposals come with two important caveats. First, the proposed codes in such low-overhead distillation schemes have high-weight parity checks with check weights that are typically extensive in the block length of the code. As such these codes are not obviously fault-tolerant on their own in the presence of circuit-level noise as they require syndrome extraction circuits with depths that grow with the block-length, and thus may not support a finite threshold or exponential sub-threshold error suppression. Morever, the extensive depth of syndrome extraction circuits of these codes implies that the temporal overhead associated with distillation from these codes is significant. Second, some of these constructions require constant-depth Clifford corrections on top of a transversal non-Clifford circuit to achieve the desired logical action. These schemes implicitly assume that Clifford operations are noiseless – an assumption that is justified when concatenating the distillation code with an inner code that has transversal Clifford operations such as the two-dimensional color code. Practically, the overhead of these protocols is therefore exaggerated by the requirement of a sufficiently high distance inner code so that the distillation protocol is not bottlenecked by the error-rate of Clifford operations.
@@
Quantum low-density parity-check (qLDPC) codes, which simultaneously admit sparse stabilizer checks and favorable rate and distance scaling, have been proposed as a pathway toward reducing the cost of quantum error correction. In the context of quantum memory, qLDPC codes with asymptotically constant rates and relative distances have recently been shown to exist [67, 53], and several architectures for implementing qLDPC memories on superconducting and neutral-atom architectures have been proposed [5, 95, 96, 13]. However, the application of general qLDPC codes to magic state distillation protocols has only recently begun to be explored, in part because of the difficulty of constructing codes that host transversal non-Clifford gates. Utilizing appropriate qLDPC codes could dramatically reduce the overhead of MSD protocols for a number of reasons. Codes with high relative distances ensure effective error suppression during distillation, while high rate codes allow multiple magic states to be distilled in parallel within a single block, improving throughput. In contrast to traditional MSD schemes that rely on small codes concatenated many times, LDPC codes can potentially achieve the same target fidelity in a single round, greatly reducing space-time overhead. While this is also true for the aforementioned non-LDPC constant-overhead and low-overhead MSD protocols, crucially, qLDPC codes enable these properties while maintaining fault-tolerance with constant-depth syndrome extraction circuits. Moreover, for qLDPC codes with strongly-transversal non-Clifford gates (i.e a constant depth physical circuit of non-Clifford gates without a Clifford correction), there is no need for concatenation with an inner code with transversal Clifford gates. Constructions of practical LDPC magic state factories are thus highly desirable for fault-tolerant quantum computing architectures.
@@
Motivated by these considerations, this work introduces a class of finite block-length, high-rate and distance quantum LDPC codes with a transversal logical $CCZ$ circuit action for magic state distillation, which we call ‘tricycle codes’. These codes are constructed as the three-dimensional balanced product [17], a particular generalization of the homological product, of classical binary linear codes over Abelian group algebras [44]. This construction extends the two-block group algebra codes of [49, 54] (also known as bicycle codes [65]) into three dimensions, allowing in principle for transversal gates from the third level of the Clifford hierarchy [37]. We may thus also refer to the tricycle codes as ‘three-block Abelian group-algebra codes’. It is worth highlighting that for the task of magic state distillation, high-rate and high-distance finite block-length codes are sufficient for essentially any large-scale quantum computation. For example, codes with distances large enough to achieve a logical error rate of $10^{-12}$ at attainable physical error rates are expected to be sufficient to produce magic states to execute large-scale quantum algorithms such as Shor’s algorithm [78] fault-tolerantly. We present several examples of finite block-length codes that achieve low logical error rates under circuit level noise at a physical error rate of $10^{-3}$.
@@
The theoretical formalism we utilize for constructing transversal non-Clifford gates in three-dimensional product codes comes from endowing the codes with a cup product, an operation from algebraic topology and homological algebra, related to the triple-intersection of logical operators of the code, which we discuss further in Sec.˜2.2 and Appendix˜B. At a high level, this formalism allows general product quantum CSS codes to be equipped with the necessary structure to host transversal gates ascending the Clifford hierarchy. In fact, the transversal non-Clifford gates of three-dimensional color codes can also be interpreted in terms of cup-products on the associated three-dimensional cellular complexes [9]. Recent work by Breuckmann et al. introduced a general formalism to endow the co-chain complexes associated with homological product and balanced product quantum codes with cup-product operations and classified the associated gates they give rise to [16]. In this work, we leverage this formalism for the tricycle codes and study the transversal logical non-Clifford circuits they host in detail. We present several concrete examples of these codes with high rates and distances, study their performance under realistic circuit level noise, and explore their application to practical magic state distillation protocols.
@@ 1.2 Summary of results
The tricycle codes we construct achieve favorable parameters at relatively small block lengths. We introduce several codes with check weights $w_{x}=12$, $w_{z}=8$, where $w_{x}$ and $w_{z}$ denote the weights of the $X$ and $Z$ checks of the code respectively. We use the notation $[[N,K,D]]$ to report the block-length $N$, number of encoded logical qubits $K$, and minimum distance $D$ of quantum codes. We find codes which encode $15$ logical qubits using $375$ physical qubits, with parameters $375,15,\leq 15]]$, as well as some $405,9,\leq 20]]$ codes. Smaller codes such as those with parameters $[[192,27,8]]$ and $[[72,9,6]]$ are also found. The largest code we study in detail is a $648,18,\leq 18]]$ code. A list of several such codes along with their parameters is presented in footnote˜2.
@@
We analyze the performance of select tricycle codes under a realistic circuit level noise model in Sec.˜2.4 using a belief-propagation decoder with order-statistics post-processing (BPOSD) [72]. By comparing the performance of codes across a range of block-lengths, we are able to ascertain a high-threshold for tricycle codes under circuit-level noise of above $0.4\%$ with BPOSD decoding. We note that although the codes we present are finite-block length codes, the construction allows us to define a family of codes by choosing a single high-performing tricycle code from a given block-length. Although such a code family is unlikely to be asymptotically good, we expect that a larger block-length code with a larger distance can always be found – a claim that is supported by the distance lower bound we prove in Appendix˜A. Defining a threshold with respect to such a constructed code family allows one to make guarantees of sub-threshold noise suppression by selecting codes of arbitrarily large block-lengths with larger distances. We expect the quoted $0.4\%$ threshold to improve significantly with better decoders, as we find that circuit level performance is highly sensitive to BPOSD hyperparameters.
@@
We also show how to construct compact syndrome extraction circuits of optimal depth (depth $12$ excluding state preparation and readout for the $w_{x}=12$, $w_{z}=8$ codes we consider). Finally, we present a concrete protocol for efficiently implementing the preparation and syndrome extraction of tricycle codes on a reconfigurable neutral-atom array platform – a video illustrating the atom-moving scheme for syndrome extraction of tricycle codes is provided in the ancillary files that accompany this manuscript.
@@
Our work serves as a step toward a new class of distillation protocols that leverage the structural properties of quantum LDPC codes to reduce resource overhead in universal fault-tolerant computation. Such distillation schemes are naturally suited to fault-tolerant quantum computing architectures that utilize LDPC quantum memories along with universal adapters for magic state injection, such as in Ref. [96, 97]. An important open problem that would make the ideas presented in this work even more efficient is to construct schemes for the efficient extraction of prepared magic states from LDPC factories in parallel, instead of teleporting gates into a computational code one logical qubit at a time.
@@ 1.3 Related work
Several recent works introduce binary quantum LDPC code families with asymptotically high rates and relative distances with transversal logical $CCZ$ action. Golowich and Lin [35] constructed a family of codes based on homological products of classical Sipser-Spielman expander codes [80] with local Reed-Muller codes [69, 62] with near-constant rate and relative distance that supports transversal logical non-Clifford gates. These codes however have check-weights that are not constant but grow logarithmically in the block-length. Lin also introduced a general formalism for constructing LDPC quantum code families with transversal non-Clifford gates based on algebraic sheaves [55]. Similarly, Zhu introduced a family of topological codes with constant rate and polynomial distance with transversal CCZ gates [100, 101]. Quantum rainbow codes were introduced as generalizations of color-codes with transversal non-Clifford gates to higher-dimensional simplicial complexes [77], achieving asymptotically constant rate and logarithmic distance. Finally, Breuckmann et al. [16] discuss how to construct constant rate and polynomial distance families with transversal gates that ascend the Clifford hierarchy using iterated homological and balanced products of chain complexes and the cup-product gate formalism they develop, which we also utilize in this work.
Adding to these remarkable results, this work focuses on a particular class of high-performance finite block-length codes. It is unclear precisely how large the smallest elements of the code families in the aforementioned works are, but constructions based on three-fold homological products of classical codes likely involve at-least several thousand and potentially tens of thousands of physical qubits for the smallest attainable block-lengths with good parameters, due to the cubic growth of the size of the quantum code with respect to the classical codes. The significantly smaller block-lengths of the tricycle codes presented in this work is one of the primary reason we chose to study them. A direct comparison of overhead and performance of the LDPC codes from the above constructions with each other and the tricycle codes we present would be an interesting direction for future work.
@@ 2.1 Tricycle codes
where $\mathbf{A}=\sum_{i=1}^{w_{a}}\mathbf{A_{i}}\in\mathbb{F}_{2}^{n_{G}\times n_{G}}$, $\mathbf{B}=\sum_{i=1}^{w_{b}}\mathbf{B_{i}}\in\mathbb{F}_{2}^{n_{G}\times n_{G}}$, $\mathbf{C}=\sum_{i=1}^{w_{c}}\mathbf{C_{i}}\in\mathbb{F}_{2}^{n_{G}\times n_{G}}$ for constants $w_{a},w_{b},w_{c}$, and $\mathbf{A_{i}},\mathbf{B_{i}},\mathbf{C_{i}}$ are permutation matrices – i.e matrices with a single 1 in every row and column. The parameter $n_{G}$ is the linear size of each block and is related to the size of a finite group (hence the subscript $G$) – the details of this connection are discussed in Appendix˜A. These parity check matrices thus define codes on $N=3n_{G}$ qubits. Modulo two arithmetic implies immediately that the CSS orthogonality condition $H_{Z}H_{X}^{T}=0$ is satisfied. Moreover, the matrices $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$ commute pair-wise. There are $n_{G}$ many $X$-checks of weight $w_{a}+w_{b}+w_{c}$ and $3n_{G}$ many $Z$-checks which can be partitioned into three sets of $n_{G}$ checks each with weights $w_{c}+w_{a}$, $w_{c}+w_{b}$, and $w_{b}+w_{a}$ respectively. By choosing $w_{a}=w_{b}=w_{c}=w$ for a constant $w$, these codes are $(w,3w)$-regular LDPC codes in the $X$ basis and $(2w,2w)$-regular LDPC in the $Z$ basis, where $(p,q)$-regular means that every qubit is involved in exactly $p$ checks and each check involves exactly $q$ qubits. Examples of particular codes from this construction in footnote˜2 demonstrate that the checks are typically highly redundant – often leading to favorable code parameters and ease of decoding – since the total number of checks $4n_{G}\gg N-K=3n_{G}-K$. This observation is encapsulated by the fact that tricycle codes possess meta-checks in the $Z$ basis – these are relations between the $Z$ checks encoded in a matrix $H_{meta}$ such that $H_{meta}\cdot H_{z}=0$. We discuss the meta-check structure and its implications in Sec.˜2.3.
@@
To ensure practical feasibility of implementing these codes in early fault-tolerant architectures below pseudo-threshold error rates, we restrict our attention to codes constructed with $w=4$, producing codes with a maximum check weight of $12$ – in particular, the $X$-checks have weight $12$ while the $Z$-checks have weight $8$. There are two reasons we consider $w=4$. The first is that the cup-product transversal $CCZ$ gate formalism of Ref. [16] that we use requires $w$ to be even for these codes (details are in Appendix˜B). The second is that we empirically found that all codes with $w=2$ had a maximum distance of $6$ and relatively poor rates regardless of the code block-length (to the extent of our numerical search).
$\mathbf{N}$ $\mathbf{K}$ $\mathbf{D_{X}}$ $\mathbf{D_{Z}}$
$72$ 99 $12$ 66
$144$ 66 $\leq 24$ $10$
$192$ $27$ $16$ 88
$375$ $15$ $\leq 25$ $\leq 15$
$405$ 99 $30$ $\leq 15$
$648$ $18$ $\leq 48$ $\leq 18$
@@
Table 1: Examples of tricycle codes and their parameters for various block-lengths. NN is the block-length, $K$ the number of encoded logical qubits, $D_{X}$ and $D_{Z}$ are the minimum weights of logical $X$-type and $Z$-type operators respectively. The distances of codes are found exactly by using either a SAT-solver to find the shortest error [32] or a mixed integer program [51] when feasible. Otherwise, distances are estimated – sometimes, with strong statistical guarantees, in which case we report the distance as exact [^2] Specifically, the method we use returns a confidence interval on the distance conditioned on certain assumptions on the distribution of minimum-weight code-words. These assumptions can in turn be hypothesis tested using the empirical distribution of low-weight code-words found from sampling. We report the distance as exact if the confidence is $99.9\%$ and if the $p$-value associated with the hypothesis test is $10\%$. Details of the method are in Appendix˜D .– using the low-weight biased Monte-Carlo sampling method discussed in Appendix˜D with 100 million shots. The construction of these specific codes is described in Appendix˜A and LABEL:table::code_polys.
@@
The codes in footnote˜2 have imbalanced distances with $d_{Z}\leq d_{X}$. Based on an extensive numerical search over several tens of millions of codes, we conjecture that this is true for all tricycle codes (Conjecture 1 in Appendix˜A). The imbalanced distances of these codes is a particularly useful feature for experimental platforms which are typically dominated by noise biased towards one basis. A detailed exploration of how noise bias can boost the experimental performance of tricycle codes is left for future work.

### v2

@@ Abstract
The preparation of high-fidelity non-Clifford (magic) states is an essential subroutine for universal quantum computation, but imposes substantial space-time overhead. Magic state factories based on high rate and distance quantum low-density parity check (LDPC) codes equipped with transversal non-Clifford gates can potentially reduce these overheads significantly, by circumventing the need for multiple rounds of distillation and by producing a large number of magic states in a single code-block. As a step towards realizing efficient, fault-tolerant magic state production, we introduce a class of finite block-length quantum LDPC codes which we name tricycle codes, generalizing the well-known bicycle codes to three homological dimensions. These codes can support constant-depth physical circuits that implement logical $CCZ$ gates between three code blocks. To construct these constant-depth $CCZ$ circuits, we develop new analytical and numerical techniques that apply to a broad class of three-dimensional homological and balanced product codes. We further show that tricycle codes enable single-shot state-preparation and error correction, leading to a highly efficient magic-state generation protocol. Numerical simulations of specific codes confirm robust performance under circuit-level noise, demonstrating a high circuit-noise threshold of $0.5\%$. With modest post-selection, certain tricycle codes of block-lengths of only 5010050-100 qubits are shown to achieve logical error-rates of $6\times 10^{-10}$ or lower. Finally, we construct optimal depth syndrome extraction circuits for tricycle codes and present a protocol for implementing them efficiently on a reconfigurable neutral atom platform.
@@ 1.1 Background
Over the last decade, remarkable progress has been made towards significantly reducing the overhead associated with magic state distillation by designing various codes with transversal non-Clifford gates and favorable code parameters [16, 54, 43, 23]. In fact, it has very recently been shown that families of quantum codes with transversal non-Clifford gates and asymptotically constant rate and relative distance exist, leading in principle to constant spatial overhead MSD protocols [73, 106]. However, these proposals come with two important caveats. First, the proposed codes in such low-overhead distillation schemes have high-weight parity checks with check weights that are typically extensive in the block length of the code. As such these codes may not be obviously fault-tolerant on their own in the presence of circuit-level noise as they require syndrome extraction circuits with depths that grow with the block-length, and thus may not support a finite threshold or exponential sub-threshold error suppression with bare-ancilla based syndrome extraction. Second, these schemes implicitly assume that Clifford operations are noiseless to execute the Clifford encoding and unencoding circuits of the codes [17] – an assumption that is justified when concatenating the distillation code with an inner code that has transversal Clifford operations such as the two-dimensional color code. Practically, the overhead of these protocols is therefore exaggerated by the requirement of a sufficiently high distance inner code so that the distillation protocol is not bottlenecked by the error-rate of Clifford operations. Moreover, the Clifford encoding and unencoding circuits can have depths scaling linearly in the number of physical qubits of the code, leading to significant temporal overhead of such protocols.
@@
Quantum low-density parity-check (qLDPC) codes, which simultaneously admit sparse stabilizer checks and favorable rate and distance scaling, have been proposed as a pathway toward reducing the cost of quantum error correction. In the context of quantum memory, qLDPC codes with asymptotically constant rates and relative distances have recently been shown to exist [78, 59], and several architectures for implementing qLDPC memories on superconducting qubit and neutral-atom architectures have been proposed [7, 107, 109, 15]. However, the application of general qLDPC codes to magic state generation protocols has only recently begun to be explored, in part because of the difficulty of constructing codes that host transversal non-Clifford gates. Utilizing appropriate qLDPC codes could dramatically reduce the overhead of MSD protocols for a number of reasons. Codes with high relative distances ensure effective error suppression during distillation, while high rate codes allow multiple magic states to be distilled in parallel within a single block, improving throughput. In contrast to traditional MSD schemes that rely on small codes concatenated many times, LDPC codes can potentially achieve the same target fidelity in a single round, greatly reducing space-time overhead. While this is also true for the aforementioned non-LDPC constant-overhead and low-overhead MSD protocols, crucially, qLDPC codes enable these properties while maintaining fault-tolerance with constant-depth syndrome extraction circuits. Moreover, for qLDPC codes with strongly-transversal non-Clifford gates (i.e a constant depth physical circuit of non-Clifford gates without a Clifford correction), there is no need for concatenation with an inner code with transversal Clifford gates. Constructions of practical LDPC magic state factories are thus highly desirable for fault-tolerant quantum computing architectures.
@@
Motivated by these considerations, this work introduces a class of finite block-length, high-rate and distance quantum LDPC codes with a transversal logical $CCZ$ circuit action for magic state distillation, which we call ‘tricycle codes’. These codes are constructed as the three-dimensional balanced product [19], a particular generalization of the homological product, of classical binary linear codes over Abelian group algebras [48]. This construction extends the two-block group algebra codes of [53, 62] (also known as bicycle codes [76]) into three dimensions, allowing in principle for transversal gates from the third level of the Clifford hierarchy [40]. We may thus also refer to the tricycle codes as ‘three-block Abelian group-algebra codes’. We highlight that for the task of magic state distillation, high-rate and high-distance finite block-length codes are sufficient for essentially any large-scale quantum computation. For example, codes with distances large enough to achieve a logical error rate of $10^{-12}$ at attainable physical error rates are expected to be sufficient to produce magic states to execute large-scale quantum algorithms such as Shor’s algorithm [89] fault-tolerantly.
@@
The theoretical formalism we utilize for constructing transversal non-Clifford gates in three-dimensional product codes comes from endowing the codes with a cup-product, an operation from algebraic topology and homological algebra, related to the triple-intersection of logical operators of the code, which we discuss further in Sec.˜2.2 and Appendix˜D. At a high level, this formalism allows general product quantum CSS codes to be equipped with the necessary structure to host transversal gates ascending the Clifford hierarchy. In fact, the transversal non-Clifford gates of three-dimensional color codes can also be interpreted in terms of cup-products on the associated three-dimensional cellular complexes [11, 55]. Recent work by Breuckmann et al. introduced a general formalism to endow the cochain complexes associated with homological product and balanced product quantum codes with cup-product operations and classified the associated gates they give rise to [18]. In this work, we first present a modification of this formalism, leading to a set of conditions that high rate and distance tricycle codes with non-trivial $CCZ$ action must satisfy (see Appendix˜D). The conditions resulting from this modified cup-product construction are essential to finding tricycle codes with both high rates and relative distances. We then study the transversal non-Clifford circuits hosted by codes which satisfy these conditions in detail. More generally, we also introduce a new numerical method that is inspired by the cup-product construction for finding short-depth code-space preserving non-Clifford circuits on iterated balanced-product quantum codes, which we also apply to certain tricycle codes. We present several concrete examples of such high rate and distance tricycle codes, study their performance under realistic circuit level noise, and explore their application to practical magic state distillation protocols.
@@ 1.2 Summary of results
The tricycle codes we construct achieve favorable parameters at relatively small block lengths, while hosting constant-depth circuits that implement a logical $CCZ$ action. We introduce several classes of tricycle codes with check weights $w_{x}\leq 12$, $w_{z}\leq 8$, where $w_{x}$ and $w_{z}$ denote the weights of the $X$ and $Z$ checks of the code respectively. We use the notation $[[N,K,D]]$ to report the block-length $N$, number of encoded logical qubits $K$, and minimum distance $D$ of quantum codes. Some examples of codes we find have parameters $[[48,6,4]]$, $[[108,21,6]]$, $[[270,24,8]]$, $[[270,12,12]]$, $480,12,\leq 17]]$, exhibiting highly competitive distances and rates that are compatible with a transversal non-Clifford action. A more complete list of codes along with the depths of the non-Clifford transversal circuits they host can be found in footnote˜2.
@@
We analyze the performance of select tricycle codes under a realistic circuit level noise model. By comparing the performance of codes across a range of block-lengths, we are able to ascertain a threshold of $0.5\%$ under circuit-level noise with an MLE decoder for a certain family of tricycle codes. We also study how post-selection and full error-detection can improve the performance of these codes. To do so, we utilize a recently developed efficient cluster-based post-selection method for quantum LDPC codes [57] combined with Belief-Propagation + Localized Statistics Decoding (BP+LSD) [46]. With full error detection, we show that even the smallest tricycle code we study in detail (with parameters $[[48,6,4]]$) can achieve logical error rates of approximately $6\times 10^{-10}$ at an acceptance fraction of roughly $30\%$. For a larger code, we estimate logical error rates $10^{-12}$ with full error detection at a $\approx 9\%$ acceptance fraction, compatible with full-scale fault tolerant quantum computation. Note that these logical error rates are achieved in a numerical simulation of the code as a memory under multiple rounds of error correction (with post-selection). In particular, they do not account for the error introduced by the $CCZ$ circuit when determining the fidelity of the final magic-state. However, since the $CCZ$ circuits are relatively short (depth $8$ for the codes discussed above), we anticipate that these estimates from the logical memory performance are a close proxy to the fidelity of the target magic-state – we discuss this perspective in more detail in Sec.˜2.4.
@@
We note that although the codes we present are finite-block length codes, the construction allows us to define a family of codes by choosing a single high-performing tricycle code from a given block-length. Although such a code family is unlikely to be asymptotically good, we expect that a larger block-length code with a larger distance can always be found – a claim that is supported by the distance lower bound we prove in Appendix˜A. Defining a threshold with respect to such a constructed code family allows one to make guarantees of sub-threshold noise suppression by selecting codes of arbitrarily large block-lengths with larger distances.
@@
We then show how to construct compact syndrome extraction circuits of optimal depth for all tricycle codes. Finally, we present a concrete protocol for efficiently implementing the preparation and syndrome extraction of these codes on a reconfigurable neutral-atom array platform – a video illustrating the atom-moving scheme for syndrome extraction of tricycle codes is provided in the ancillary files that accompany this manuscript.
We also clarify some subtle points about the application of the cup-product gate formalism of [18]. In particular, the success of this formalism for producing transversal non-Clifford circuits on tricycle codes with high rate and distance requires a generalization of the conditions presented in section $5$ of [18]. These resulting conditions will likely also be useful for other homological product and balanced product code families. We discuss these points on the theory of cup-product based gates in Appendix˜D. In Appendix˜E, we also present a complementary numerical method that generalizes the cup-product construction for finding code-space preserving $CCZ$ circuits on three-dimensional balanced-product codes.
@@
Our work serves as a step toward a new class of distillation protocols that leverage the structural properties of quantum LDPC codes to reduce resource overhead in universal fault-tolerant computation. Such magic-state generation schemes are naturally suited to fault-tolerant quantum computing architectures that utilize LDPC quantum memories along with universal adapters for magic state injection, such as in Ref. [109, 110] or using efficient code-switching protocols with lower-dimensional product codes [60, 101, 37].
@@ 1.3 Related work
Several recent works introduce binary quantum LDPC code families with asymptotically high rates and relative distances with transversal logical $CCZ$ action. Golowich and Lin [38] constructed a family of codes based on homological products of classical Sipser-Spielman expander codes [91] with local Reed-Muller codes [80, 71] with near-constant rate and relative distance that supports transversal logical non-Clifford gates. These codes however have check-weights that are not constant but grow logarithmically in the block-length. Lin also introduced a general formalism for constructing LDPC quantum code families with transversal non-Clifford gates based on algebraic sheaves [63]. Similarly, Zhu introduced a family of topological codes with constant rate and polynomial distance with transversal $CCZ$ gates [114, 115]. Quantum rainbow codes were introduced as generalizations of color-codes with transversal non-Clifford gates to higher-dimensional simplicial complexes [88], achieving asymptotically constant rate and logarithmic distance. Finally, Breuckmann et al. [18] discuss how to construct constant rate and polynomial distance families with transversal gates that ascend the Clifford hierarchy using iterated homological and balanced products of chain complexes and the cup-product gate formalism they develop.
Adding to these results, this work focuses on a particular class of high-performance finite block-length codes. It is unclear precisely how large the smallest elements of the code families in the aforementioned works are, but constructions based on three-fold homological products of classical codes likely involve at-least several thousand and potentially tens of thousands of physical qubits for the smallest attainable block-lengths with good parameters, due to the cubic growth of the size of the quantum code with respect to the classical codes. The significantly smaller block-lengths of the tricycle codes presented in this work is one of the primary reasons we chose to study them. A direct comparison of the overhead and performance of the LDPC codes from the above constructions with each other and the tricycle codes we present would be an interesting direction for future work.
@@ 2.1 Tricycle codes
where $\mathbf{A}=\sum_{i=1}^{w_{a}}\mathbf{A_{i}}\in\mathbb{F}_{2}^{n_{G}\times n_{G}}$, $\mathbf{B}=\sum_{i=1}^{w_{b}}\mathbf{B_{i}}\in\mathbb{F}_{2}^{n_{G}\times n_{G}}$, $\mathbf{C}=\sum_{i=1}^{w_{c}}\mathbf{C_{i}}\in\mathbb{F}_{2}^{n_{G}\times n_{G}}$ for constants $w_{a},w_{b},w_{c}$, and $\mathbf{A_{i}},\mathbf{B_{i}},\mathbf{C_{i}}$ are permutation matrices – i.e matrices with a single 1 in every row and column. The parameter $n_{G}$ is the linear size of each block and is related to the size of a finite group (hence the subscript $G$) – the details of this connection are discussed in Appendix˜A. These parity check matrices thus define codes on $N=3n_{G}$ qubits. Modulo two arithmetic implies immediately that the CSS orthogonality condition $H_{Z}H_{X}^{T}=0$ is satisfied. Moreover, the matrices $\mathbf{A}$, $\mathbf{B}$, $\mathbf{C}$ commute pair-wise. There are $n_{G}$ many $X$-checks of weight $w_{a}+w_{b}+w_{c}$ and $3n_{G}$ many $Z$-checks which can be partitioned into three sets of $n_{G}$ checks each with weights $w_{c}+w_{a}$, $w_{c}+w_{b}$, and $w_{b}+w_{a}$ respectively. By choosing $w_{a},w_{b},w_{c}\leq w$ for a constant $w$, the resulting codes are $(\leq w,\leq 3w)$-LDPC codes in the $X$ basis and $(\leq 2w,\leq 2w)$-LDPC in the $Z$ basis, where $(p,q)$-LDPC means that every qubit is involved in at most $p$ checks and each check involves at most $q$ qubits. Examples of particular tricycle codes in footnote˜2 demonstrate that the checks are typically highly redundant – often leading to favorable code parameters and ease of decoding – since the total number of checks $4n_{G}\gg N-K=3n_{G}-K$. This observation is encapsulated by the fact that tricycle codes possess metachecks in the $Z$ basis – these are relations between the $Z$ checks encoded in a matrix $H_{meta}$ such that $H_{meta}\cdot H_{z}=0$. We discuss the metacheck structure and its implications in Sec.˜2.3.
@@
To ensure practical feasibility of implementing these codes in early fault-tolerant architectures below pseudo-threshold error rates, we restrict our attention to tricycle codes with a maximum check weight at most $12$. More specifically, we consider three types of tricycle codes. The first is constructed with $w_{a}=4$, $w_{b}=w_{c}=2$, producing codes with weight 88 $X$-checks, and $Z$-checks of weight $4$ and $6$ – we call these codes $\mathbf{4-2-2}$ codes, referring to the weights $w_{a},w_{b},w_{c}$. The second class of codes we consider has $w_{a}=w_{b}=4$, $w_{c}=2$, and therefore has $X$-checks of weight $10$ and $Z$-checks of weight $8$ and $6$ – we refer to these as $\mathbf{4-4-2}$ codes. Finally, we consider codes with $w_{a}=w_{b}=w_{c}=4$, for which the $X$-checks have weight $12$ and the $Z$-checks have weight $8$ – we call these $\mathbf{4-4-4}$ codes. One may similarly consider $\mathbf{2-2-2}$ codes with weight 66 $X$-checks and weight 44 $Z$-checks, but we empirically found that the rates and distances of such codes with a non-trivial transversal $CCZ$ action are not as favorable. In particular, we could not find any such codes with $K>3$ encoded logical qubits and distance $D>7$ – such 2222-2-2 codes with $K=3$ encoded logical qubits and distance $D\leq 7$ have been studied in other recent works [51, 60].
The reason we consider $w_{a},w_{b},w_{c}$ to be $2$ or $4$ is that the specific conditions for tricycle codes to host transversal $CCZ$ circuits (see Appendix˜D) cannot be satisfied when any of $w_{a,b,c}=3$, and $w_{a,b,c}=1$ leads to a trivial code. One may consider $w_{a,b,c}>4$ to search for codes with even more favorable parameters than the ones we present here, but we empirically found 6226-2-2 codes to have parameters no better than 4224-2-2 and 4424-4-2 codes, while going to even higher check-weight will likely lead to poor thresholds.
$\mathbf{N}$ $\mathbf{K}$ $\mathbf{D_{X}}$ $\mathbf{D_{Z}}$ Type $\mathbf{CCZ}$ depth $\mathbf{CCZ}$ method
$48$ 66 88 44 4224-2-2 88 STCP
$84$ 66 $12$ 55 4224-2-2 88 STCP
$108$ 66 $12$ 66 4224-2-2 88 STCP
$240$ 66 $\leq 22$ 88 4224-2-2 88 STCP
$480$ 66 $\leq 36$ $10$ 4224-2-2 88 STCP
$108$ $12$ $11$ 44 4424-4-2 $16$ STCP
$180$ $12$ $15$ 66 4424-4-2 $32$ STCP
$108$ $15$ $12$ 66 4444-4-4 $96$ NLR
$270$ $24$ $15$ 88 4444-4-4 $96$ NLR
$324$ $12$ $\leq 32$ $\leq 12$ 4444-4-4 $128$ STCP
$480$ $15$ $\leq 48$ $\leq 14$ 4444-4-4 $128$ STCP
@@
Table 1: Examples of tricycle codes and their parameters for various block-lengths. NN is the block-length, $K$ the number of encoded logical qubits, $D_{X}$ and $D_{Z}$ are the minimum weights of logical $X$-type and $Z$-type operators respectively. The distances of codes are found exactly by using either a SAT-solver to find the shortest error [35] or a mixed integer program [56] when feasible. Otherwise, distances are estimated – sometimes, with strong statistical guarantees, in which case we report the distance as exact [^2] Specifically, the method we use returns a confidence interval on the distance conditioned on certain assumptions on the distribution of minimum-weight code-words. These assumptions can in turn be hypothesis tested using the empirical distribution of low-weight code-words found from sampling. We report the distance as exact if the confidence is $99.9\%$ and if the $p$-value associated with the hypothesis test is $10\%$. Details of the method are in Appendix˜C.– using the low-weight biased Monte-Carlo sampling method discussed in Appendix˜C with 100 million shots. The construction of these specific codes is described in Appendix˜A and LABEL:table::code_polys. The column named ‘Type’ denotes the class of tricycle codes, labeled by the weights $w_{a}-w_{b}-w_{c}$ from Eq.˜1. The depths of the code-space preserving $CCZ$ circuits with non-trivial logical action are noted for each code. The final column named ‘CCZCCZ Method’ indicates whether the code and non-trivial CCZCCZ circuit was constructed using the symmetric triple cup-product method (STCP) – see Appendix˜D – or using the numerical Leibniz rule method (NLR) – see Appendix˜E.
@@
The codes in footnote˜2 have imbalanced distances with $d_{Z}\leq d_{X}$. In Appendix˜A, we prove that this is the case for all tricycle codes. The imbalanced distances of these codes is a potentially useful feature for experimental platforms which are typically dominated by noise biased towards one basis. A detailed exploration of how noise bias can boost the experimental performance of tricycle codes is left for future work.
@@
For the remainder of this work, we primarily focus on select 4224-2-2 type tricycle codes that are practically relevant due to their lower check-weights and depth-$8$ CCZCCZ circuits. The 4444-4-4 tricycle codes generally exhibit higher rates and distances, but this comes at the price of higher check-weights and a lower circuit-noise threshold of $\approx 0.4\%$, as well as significantly deeper non-trivial $CCZ$ circuits on average – additional details on the 4444-4-4 codes are presented in Appendix˜D, Appendix˜E, and Appendix˜G. We empirically found that the 4424-4-2 type codes did not offer substantial improvements in rates, distances, or thresholds compared to the 4224-2-2 and 4444-4-4 type codes to the extent of our code search – two examples of 4424-4-2 codes with non-trivial $CCZ$ are nonetheless presented in footnote˜2.